How do you solve this problem using the substitution method?
5x+4y=7
y=-3x

Answer by jim_thompson5910(35256)

since y = -3x, we can replace the 'y' in the first equation with '-3x'

so we go from

5x+4y=7

to

5x + 4(-3x) = 7

then multiply to get

5x - 12x = 7

Keep going to solve for x. Once you've found x, plug it into y = -3x to find y.